6) A ball is thrown in a projectile motion trajectory with an initial velocity v at an angle θ above the ground. If the acceleration due to gravity is which of the following is the correct expression of the time it takes for the ball to reach its highest point, y, from the ground? a. vsin c. vsie
Answer #1

Solution:

The horizontal component of the velocity is v cos(θ) and vertical component is vsin(θ)

Now Using the formula v = u+at in vertical direction

=> 0 = vsin(θ) - g * t

=> t = vsin(θ) / g

Hence option c should be right.

thanks

please rate.